# Diffusion as Reasoning: Enhancing Object Goal Navigation with LLM-Biased Diffusion Model

### v1

@@ III-C Semantic Reasoning with Local Map Condition
The observed local semantic map can be used as a condition to guide the diffusion model in generating the object distribution in the unknown regions, thereby enabling semantic reasoning of the target objects. The local map established for the explored regions is denoted as ${x}$. After cropping and scaling preprocessing, the unknown area in ${x}$ constitute the mask ${m}$. The known grid is represented as ${x\odot\left(1-m\right)}$, and the corresponding unknown region is denoted as ${x\odot m}$. We apply the RePaint [39] method, using Eq. 5 for the known region and Eq. 6 for the unknown region.
@@
${x_{t-1}^{unkown}\sim\mathcal{N}\left(\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t\right)\right)}$ (6)
As shown in Figure 5, ${\bm{x}_{\bm{t}-\bm{1}}}$ and ${\bm{x}_{\bm{t}-\bm{1}}^{\bm{\theta}}}$ represent the sampled result from the local map and the model trained in the previous section, respectively.
@@
They are combined through the process represented by Eq. 7 and participate in the reverse process, ultimately generating a reasonable semantic content distribution in the unknown region of the local map.
III-D Enhancing Generation with LLM Bias
Although we found that the diffusion model can effectively generate the object distribution in unknown regions, the object goal navigation domain lacks large-scale datasets, making it difficult to achieve satisfactory generalization through supervised training. Recently, the ability of knowledge extraction and integration within Large Language Models (LLMs) for robotics has been extensively investigated. LLMs can provide agents with common sense knowledge about the environment, which can improve the success rate of object goal navigation tasks [22][19].
The research in Section III-B showed that by changing the channel bias of the initial noise in the denoising model for map generation, the richness of each object in the final semantic map can be controlled. This bias is global. In contrast, this section proposes a local bias method that uses LLMs to predict the object distribution in different regions and applies specific biases in those regions, thereby exerting control over the diffusion model’s map generation process.
Figure 4: Prompts and object association matrix.
We collect the common sense knowledge by querying the LLM with positive-negative prompts [22] and Chain-of-Thought (CoT) prompting technology [43]. Additionally, we query the LLM from the dimensions of semantic and spatial relevance, and then fuse the statistical results of these two dimensions to generate the common sense knowledge matrix ${M_{CS}}$. Subsequently, we set the diagonal values of ${M_{CS}}$ to 0.
@@ III-D Enhancing Generation with LLM Bias
Our goal is to use ${M_{CS}}$ to calculate the object distribution tendency for a specific frontier point in the local map. The frontier points are defined as the edges between explored free-spaces and unexplored areas on the local semantic map [12]. Specifically, based on the coordinates of the frontier point ${f_{p}}$, we calculate the normalized distance score between all objects in the local map ${{f_{s}=\left[s_{0},s_{1},\cdots,s_{n}\right]^{T}}}$. The score ${s_{i}}$ represents the distance score of object ${i}$ relative to ${f_{p}}$, and the closer the distance, the closer the score is to 1. Taking the Gibson dataset [3] as an example, ${n=14}$, and the score for objects not present in the map is 0.
Figure 6: The proposed local LLM bias method.
Next, we compute ${M_{CS}\times f_{s}}$ by matching the object proximity sequence corresponding to the frontier point in the local map with each row in ${M_{CS}}$. The object with the highest overall score (${{\max\left(M_{CS}\times f_{s}\right)}}$) is taken as the prediction result. If the diagonal values of ${M_{CS}}$ are 1, the prediction results will tend to generate the objects closest to the frontier point ${f_{p}}$, which, for instances, would lead to generating more chair near around chair. Therefore, we set the diagonal values to zero to reduce the weight of the nearest objects, allowing the model to predict a greater variety of objects in the frontier regions.
For example, in Figure 6, the three frontier points (denoted as $F$) are predicted by ${M_{CS}}$ to most likely contain book. The ${100\times 100}$ region filled around $F$ constitutes the frontier zone. When generating the initial noise for the diffusion model, we apply a bias in the book channel within the frontier zone, while the regions outside the frontier zone and the other object channels still use the unbiased ${\mathcal{N}\left(0,\mathbf{I}\right)}$ noise.
Figure 6-a shows the results of applying global bias, where noise is applied globally in the book channel. In contrast, Figure 6-b only applies noise bias within the frontier zone, which constrains the model to generate more of the target objects in the biased region.
III-E Object Goal Navigation with DAR
@@
The trained diffusion model only allows for the prediction of map data at a resolution of ${256\times 256}$. Therefore, we crop the minimum bounding box of the known region from the input ${m_{t}}$, as shown by the blue dashed box in Figure 7. We then rescale the box region to a size of ${256/\epsilon}$ and pad it with zeros to ${256\times 256}$, generating the processed map ${m_{t}^{p}}$. The value of ${\epsilon}$ represents the range of unknown regions that the DAR model is used to predict. We set ${{\epsilon}=120\%}$.
Figure 7: Workflow of the proposed DAR method during navigation.
Next, we generate the initial noise with biases for the sampling process based on ${m_{t}^{p}}$. First, we apply the global target bias introduced in Section III-A. Then computing the frontier points ${F_{i}}$ of ${m_{t}^{p}}$, and utilize the common sense knowledge matrix ${M_{CS}}$ extracted from the LLM, as described in Section III-C, to estimate the most likely objects for each ${F_{i}}$. Finally, we apply the local bias.
The biased noise goes through the denoising process described in Section III-B, gradually producing a semantic map ${m_{t}^{d}}$ within the unknown area of the ${m_{t}^{p}}$. In the ${m_{t}^{d}}$, non-zero values in the target channel indicate the predicted potential location of the target object by the DAR model. These coordinates are restored to the local map ${m_{t}}$ through inverse resizing and used as the long-term goal ${g_{t}}$.
It is worth noting that in the initial stage of navigation or when the sensors receive limited information, even with the introduction of global target bias, the map generated by the DAR model in the unknown regions may not contain the target object. In other words, the DAR model may not provide valuable information to the navigation system. In such cases, we set the long-term goal to the nearest frontier point.
@@
While the diffusion model can effectively predict the object distribution in unknown regions, its nature also makes the DAR method computationally expensive. To address this drawback, we do not call the DAR method to reason about the long-term goal at every timestamp. Instead, we adopt a strategy where we only perform the reasoning when the current location is near the ${g_{t}}$, or when the agent is moving away from ${g_{t}}$. Subsequent experiments have shown the effectiveness of this approach, as the DAR model can still provide a significant performance boost to the object goal navigation system, even with fewer reasoning steps.
IV Experiment
IV-A Experimental setup
Dataset. We evaluate our DAR on standard ObjectNav datasets: Gibson [3], Matterport3D (MP3D) [1], and Habitat-Matterport3D (HM3D) [23]. For Gibson and MP3D, we follow the setup of [12] and [13]. Specifically, for Gibson, we use 25 training and 5 validation scenes from the Gibson tiny split. For MP3D, we use 56 training and 11 validation scenes, with 21 goal categories and 2,195 validation episodes. For HM3D, we follow [44], using 80 training and 20 validation scenes, with 6 goal categories and 2,000 validation episodes.
@@ IV-A Experimental setup
Implementation details. For the self-supervised diffusion learning of DAR, we take room semantic maps from the Gibson and MP3D datasets, randomly rotate, scale, and crop them to 256x256 size for training samples. We use a DDPM model based on UNet, with 2 BigGAN residual blocks.
IV-B Comparisons with previous map completion methods.
@@
Figure 8: The DAR model reconstructing the map of the masked region.
As shown in Table I, we compare the proposed DAR method (a mask-agnostic approach) with the Unet method and SGM (which utilizes ViT and learns with randomly masked sub-maps). In the experiments, as illustrated in Figure 8, we use a 75% random mask consistent with SGM, using the known areas as conditions to generate the masked region maps. To examine the contribution of ${M_{CS}}$ to object prediction, we set up ${16\times 16}$ patches and randomly mask 25% of the map. For each masked patch, we use the method in Section III-D to call ${M_{CS}}$ to predict the possible objects, and then calculate the recall of the semantic map based solely on ${M_{CS}}$ predictions. The ${M_{CS}}$ extracted from the LLM does not contain any prior information about the rooms in the Gibson or other datasets, so it requires a sufficient amount of known areas to make effective judgments. Therefore, we only mask 25% of the regions in the experiments in Table II.
As shown in Table II, using the COT and positive-negative prompting technology to extract the ${M_{CS}}$ matrix representing the associative relationships between objects from the LLM can effectively improve the performance of the DAR model in navigation.
Table I: Comparisons of different map generation method
ID Model LLM RS Map generation Navigation
IoU (%$\uparrow$ Recall (%$\uparrow$ SR (%$\uparrow$ SPL (%$\uparrow$
1 UNet - - 19.28 47.36 67.5 37.8
SGM (ViT-B) - - 25.45 79.23 74.1 39.8
SGM (ViT-B) GPT-4 - 39.47 87.33 77.4 43.8
@@
DAR (0.5K steps) - 22.11 73.12 73.2 38.7
DAR (0.2K steps) - 22.31 72.96 69.5 39.3
DAR (0.2K steps) - 17.84 65.44 65.6 36.1
@@
IV-C Diffusion process analysis.
The proposed DAR model uses DDPM to reason about the object distribution in unknown regions by pre-training the DDPM on the semantic map dataset. We set up various diffusion steps and obtain the training loss decay curves shown in Figure 9. It can be seen that the higher the diffusion steps, the faster the model’s training loss decays and the lower the final convergence level, indicating that increasing the diffusion steps helps the model learn and generate more reasonable room semantic maps. This can also be observed from the second and third groups (id=2 and 3) in Table I. However, more diffusion steps during sampling means more computational time.
The REPAINT method introduces a resampling strategy to harmonize the model’s input, producing higher-quality images. We apply this approach to enhance the realism of the generated maps. The ’RS’ in Table I represents whether the resampling process is used. We find that a 0.5k steps pre-trained model with resampling can achieve higher map generation and navigation performance compared to a 1k steps model without resampling.
Figure 9: Training loss curves for different diffusion steps.
IV-D Visualization of navigating with DAR.
@@
We follow the SGM configuration, and at the beginning of the navigation stage, the agent’s collected environmental information is insufficient to support accurate visual reasoning, so we use the original exploration strategy. As more scenes are observed, the DAR model successfully predicts the position of the target (toilet), even though it has not yet been observed. During the navigation process, DAR continuously predicts the object distribution in the unknown area based on the agent’s accumulated environmental memory, and provides a long-term goal to guide the exploration direction. It is worth noting that due to the inherent characteristics of the diffusion model, DAR’s semantic reasoning contains a certain degree of randomness. Even in the same scene, with the same starting position and target, each call to DAR for navigation will yield different results.
@@
As shown in Figure 10, although the paths obtained are not completely identical, overall, DAR still significantly improves the efficiency of ObjectNav. From Figure 11, it can also be seen that the DAR model not only successfully predicts the potential location of the target object during the navigation process, but also estimates the approximate locations of other objects and their context. There is still some deviation between the predicted target position and the true position by DAR, but we believe this is acceptable, as even humans cannot precisely locate unseen objects in an unseen scene.
@@
These two visualizations show that our DAR effectively expands limited local maps by generating unobserved regions, enabling the agent to infer the target’s location.
Figure 10: Visualization of navigation using the proposed DAR method.
Figure 11: Generated maps during navigation.
IV-E Ablation study.
Table II: The impact of prompting techniques on performance
Settings Map Prediction Navigation
COT Positive Negative Relation Recall (%$\uparrow$ SR (%$\uparrow$ SPL (%$\uparrow$
both 73.6 76.1 42.8
both 74.4 77.9 43.7
both 75.0 76.4 43.1
@@
spa 75.2 76.6 43.4
both 76.7 78.3 44.2
Table III: Ablation study.
Settings Gibson
GT GTB LLB EXB SR(%$\uparrow$ SPL(%$\uparrow$ DTS(m)$\downarrow$
91.7 72.3 0.38
72.4 40.9 1.51
75.1 42.4 1.45
@@
We verify the global target bias and local LLM bias methods in the DAR model, and obtain the results shown in Table III. Introducing an exploration strategy early in navigation is effective, as it prevents being misled by low-confidence predictions from limited observations. The performance of the Local LLM bias largely depends on COT and positive-negative prompting techniques. The experiments we conducted in Table III demonstrate the effectiveness of extracting the association matrix ${M_{CS}}$ from LLM by considering both spatial and semantic relations.
IV-F Comparisons with the related works.
Table IV: Performance comparison with mainstream navigation methods.
Dataset Metrics/Method End-to-End Modular-based
DD-PPO Red-Rabbit THDA Habitat-web RIM FBE ANS SemExp PONI STUBBORN SGM DAR
Gibson SR (%$\uparrow$ 15.0 - - - - 48.5 67.1 71.1 73.6 - 78.0 78.3
SPL (%$\uparrow$ 10.7 - - - - 28.9 34.9 39.6 41.0 - 44.0 44.2
DTS (m)$\downarrow$ 3.24 - - - - 2.56 1.66 1.39 1.25 - 1.11 1.08
@@
SPL (%$\uparrow$ 1.8 7.9 11.0 10.2 17.0 10.6 9.2 10.9 12.1 13.5 14.7 15.3
DTS (m)$\downarrow$ 6.94 - 5.62 - - 5.00 6.31 6.06 5.63 5.01 4.93 5.02
We categorize existing ObjectNav works, evaluated on either the Gibson or MP3D datasets, into two approaches: 1) End-to-End: Directly outputs actions, primarily using reinforcement learning. 2) Modular-Based: Predicts a map-based long-term goal. These are listed in Table IV.
@@
  • THDA: Introduces ”Treasure Hunt Data Augmentation” to improve rewards and adaptability in new scenes.
@@
In this study, although our DAR model employs a self-supervised approach, it achieves performance comparable to existing supervised methods across all metrics on various datasets. Notably, on the Gibson dataset, our DAR model surpasses the previous state-of-the-art self-supervised method by 3.5% in Success Rate (SR), 1.9% in Success weighted by Path Length (SPL), and reduces the Distance to Success (DTS) by 0.05 meters. It is important to note that a lower DTS value indicates better performance.
V Conclusions
In this work, we propose a modular-based self-supervised method for the object goal navigation task, named DAR. We demonstrate the effectiveness of applying diffusion models to visual semantic reasoning. We first trained Denoising Diffusion Probabilistic Models based on the indoor semantic map dataset, and successfully generated high-quality and reasonable new map samples, indicating that the trained DDPM can learn the statistical distribution patterns of the spatial relationships between indoor objects, and we point out that this pattern is highly beneficial for improving the success rate and efficiency of the agent in the ObjectNav task. Subsequently, we use COT and positive-negative prompting techniques to query LLMs to obtain common sense knowledge, and incorporate it into the diffusion model’s reasoning process through the proposed local bias method. In summary, our DAR method explicitly leverages the structured knowledge from the knowledge graph extracted from LLM and implicitly utilizes the unstructured knowledge from the pre-trained diffusion model. A series of experiments have shown that the DAR method achieves comparable performance to previous methods.

### v2

@@ 3.3.2 Map-Completion via Local Map-Conditioned Generation
The observed local semantic map can be used as a condition to guide the diffusion model in generating the object distribution in the unknown regions, thereby enabling semantic reasoning of the target objects. The local map established for the explored regions is denoted as ${x}$. After cropping and scaling preprocessing, the unknown area in ${x}$ constitute the mask ${m}$. The known grid is represented as ${x\odot\left(1-m\right)}$, and the corresponding unknown region is denoted as ${x\odot m}$. We apply the RePaint [27] method, using Eq. 10 for the known region and Eq. 11 for the unknown region. They are combined through the process represented by Eq. 12 and participate in the reverse process, ultimately generating a reasonable semantic content distribution in the unknown region of the local map.
@@
${x_{t-1}^{unkown}\sim\mathcal{N}\left(\mu_{\theta}\left(x_{t},t\right),\Sigma_{\theta}\left(x_{t},t,r\right)\right)}$ (11)
@@
Algorithm 1 Map completion
1:Input: Total diffusion steps $T$, the Room Guidance $r$, trained denoising network $\psi_{\theta}\left(\cdot\right)$, local map $x_{local}$, map mask $m$.
2:Output: The completed map $m^{d}$
3:Generate sub-sequence of length $S$: $\left[\tau_{1},...,\tau_{S}\right]$
4:Sample $x_{\tau_{S}}\sim\mathcal{N}\left(0,\mathbf{I}\right)$
5:for $t$ in $\left\{\tau_{S},\tau_{S-1},...,\tau_{1}\right\}$ do
6:     $\beta_{t}=\frac{10^{-4}\left(T-t\right)+2\times 10^{-2}\left(t-1\right)}{T-1}$
7:     $\alpha_{t}=1-\beta_{t}$
8:     $\bar{\alpha}_{t}=\prod\nolimits_{n=1}^{t}{\alpha_{n}}$
9:     $\epsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)\,\,\mathrm{if}\ t>1,\mathrm{else}\ \epsilon=0$
10:     $x_{t-1}^{\mathrm{known}}=\sqrt{\bar{\alpha}_{t}}x_{local}+\left(1-\bar{\alpha}_{t}\right)\epsilon$
11:     Estimate $x_{0}=\frac{x_{t}-\psi_{\theta}\left(x_{t},t,r\right)\cdot\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$
12:     Predict $\epsilon_{t}=\frac{x_{t}-\sqrt{\bar{\alpha}_{t}}x_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$
13:     $x_{t-1}^{\mathrm{unknown}}=\sqrt{\bar{\alpha}_{t-1}}x_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t}$
14:     $x_{t-1}=m\odot x_{t-1}^{\mathrm{known}}+\left(1-m\right)\odot x_{t-1}^{\mathrm{unknown}}$
15:end for
16:Return $x_{0}$
As shown in Fig. 7, ${\boldsymbol{x}_{\boldsymbol{t}-\boldsymbol{1}}}$ and ${\boldsymbol{x}_{\boldsymbol{t}-\boldsymbol{1}}^{\boldsymbol{\theta}}}$ represent the sampled result from the local map and the model trained in the previous section, respectively.
3.4 Object Goal Navigation with DAR
@@
The trained diffusion model only allows for the prediction of map data at a resolution of ${256\times 256}$. Therefore, we crop the minimum bounding box of the known region from the input ${m_{t}}$, as shown by the blue dashed box in Fig 8. We then rescale the box region to a size of ${256/\epsilon}$ and pad it with zeros to ${256\times 256}$, generating the processed map ${m_{t}^{p}}$. The value of ${\epsilon}$ represents the range of unknown regions that the DAR model is used to predict. We set ${{\epsilon}=120\%}$.
Figure 8: During navigation, the agent builds a local semantic map ${m_{t}}$ from RGBD observations, which is processed through cropping, resizing, and padding to produce $m^{p}_{t}$. The DAR model then generates a semantic map for the unknown areas in $m^{p}_{t}$, applying our proposed room guidance generation methods to manage noise in the denoising process. The long-term goal is set based on the generated map and the target.
Next, using Eq. 8 and 9, we estimate the room category for each frontier point based on pre-extracted object-room relations from the LLM, then generate room guidance using Eq. 1.
Through the process described in Section 3.2, the model progressively denoises the data and, under the guidance of room semantics, ultimately generates a semantic map ${m_{t}^{d}}$ within the unknown area of the ${m_{t}^{p}}$. In the ${m_{t}^{d}}$, non-zero values in the target channel indicate the predicted potential location of the target object by the DAR model. These coordinates are restored to the local map ${m_{t}}$ through inverse resizing and used as the long-term goal ${g_{t}}$.
It is worth noting that in the initial stage of navigation or when the sensors receive limited information, the map generated by the DAR model in the unknown regions may not contain the target object. In other words, the DAR model may not provide valuable information to the navigation system. In such cases, we set the long-term goal to the nearest frontier point.
@@
Although diffusion models can effectively predict object distributions in unknown regions, their inherent nature results in slow DAR method inference. To address this limitation, we propose two parallel strategies to enhance map generation efficiency. First, inspired by DDIM [40], we adopt an accelerated sampling method (Alg. 1), downsampling the original $T$ diffusion steps to a subsequence of $S$ steps ($S<T$) to improve sampling speed. Second, instead of invoking DAR for long-term goal inference at every timestep, we employ a selective strategy: reasoning is triggered only when the agent reaches the vicinity of the long-term goal ($g_{t}$) or deviates significantly from it. Experimental results demonstrate the effectiveness of our approach—even with reduced inference frequency, the DAR model achieves leading performance improvements for object navigation systems.
4 Experiment
4.1 Experimental setup
Dataset. We evaluate our DAR on standard ObjectNav datasets: Gibson [46] and Matterport3D (MP3D) [4]. For Gibson and MP3D, we follow the setup of [32] and [54]. Specifically, for Gibson, we use 25 training and 5 validation scenes from the Gibson tiny split. For MP3D, we use 56 training and 11 validation scenes, with 21 goal categories and 2,195 validation episodes.
@@ 4.1 Experimental setup
Implementation details. For the self-supervised diffusion learning of DAR, we take room semantic maps from the Gibson and MP3D datasets, randomly rotate, scale, and crop them to $256\times 256$ size for training samples. For training and evaluation, we use one NVIDIA RTX4090 GPU with 24GB memory.
4.2 Comparisons with previous map completion methods.
4.2.1 Map Completion Test DataSet
In this section, we construct a map completion test set that is independent of the navigation process and solely designed to evaluate the quality of algorithms in generating maps for unknown regions. Taking Gibson as an example, we first obtain the semantic object map of a scene. Random translation and rotation are then applied as data augmentation, followed by cropping to a size of $256\times 256$. We then randomly select two locations in the navigable space of the complete map and compute the shortest path using the deterministic Fast Marching Method (FMM) [38] path planner. For each location along this path, we sample an $S\times S$ (50 grids) square patch around it to represent the explored area. These patches are aggregated to form the local map $m^{d}$, while the corresponding map mask is also generated, where known and unknown regions are denoted by 1 and 0, respectively. The objective of the map completion algorithm is to generate the semantic map in unknown regions based on the map mask.
Figure 9: Generated maps during navigation. As shown by the blue circles, the DAR model accurately predicts plausible positions of unobserved objects in unobserved areas.
To evaluate the quality of the generated maps, we employ two complementary metrics: the Intersection over Union (IoU), which measures per-pixel agreement across all semantic channels including both occupancy and object categories, and Category Recall, which quantifies the algorithm’s ability to identify all relevant objects by computing the ratio of correctly predicted object categories to the total number of ground-truth categories. These metrics collectively assess both spatial accuracy and semantic recognition performance.
4.2.2 Comparison results
@@
Figure 10: Examples from the Map Completion Test Dataset, with corresponding room guidance and map completion results.
We evaluated the map completion performance of different models on the Map Completion Test Dataset, which contains data more reflective of real-world navigation scenarios. While we consider MAE appropriate for model training, it is unsuitable for validating map completion performance because, in actual navigation, known regions are interconnected rather than widely scattered and disjointed.
As shown in Table. 1, the Diffusion model significantly outperforms the ViT and CNN model. We attribute this to their inherent mechanisms. In the MAE+ViT training approach, the map is first divided into patches of a fixed size (e.g., 16 in SGM) to distinguish known and unknown regions. Due to computational and hardware constraints, the patch size cannot be set too small, leading to known-region patches containing some unknown pixels, which increases uncertainty in map generation. Moreover, the SGM method enhances completion performance by leveraging an LLM to provide general knowledge for reasoning about contextual relationships and predicting probable objects.
In this section, we also evaluated different implementations of Room-Guidance. Both the fixed $\sigma$ approach (in Eq. 1) and the single-point method underperformed compared to the adaptive approach. As shown in Fig. 10, our DAR model successfully generates high-quality semantic maps in unknown regions. The object distribution aligns with room layouts and exhibits similarity to ground truth, demonstrating that the diffusion model can learn the distribution patterns of room objects and be effectively utilized for long-term goal reasoning in navigation.
Table 1: Performance comparison of different models on the Map Completion Test Dataset. (Highest scores are highlighted in dark blue, second-highest in light blue.)
Model Gibson MP3D
@@
Diff-M2+RG ($\sigma=5$) 0.2365 0.2647 0.1879 0.3341
Diff-M3+RG ($\sigma=5$) 0.2181 0.2606 0.1840 0.3217
@@
Diff-M1+RG (self-adaptive) 0.2728 0.2808 0.2276 0.3603
4.3 Diffusion process analysis.
In this section, we conduct a detailed investigation into how the diffusion steps affects model performance. Using the ‘Diff-M1+RG (self-adaptive)’ configuration from Table. 1 and following the methodology in Alg. 1, we vary the number of diffusion steps. As shown in Fig. 11, the generation time per sample exhibits an approximately linear relationship with diffusion steps.
Our experiments reveal a trade-off: reducing diffusion steps improves computational efficiency but leads to lower IoU in map generation and decreased navigation success rate (SR). Notably, when decreasing steps from 1000 to 500, we observe only a 2-3% degradation in both IoU and SR metrics, while achieving a 49% reduction in computation time. This suggests that 500 diffusion steps represent an optimal balance between computational efficiency and model performance. In our implementation (Alg. 1), we maintain $S=500$ and $T=1000$, the model is trained with 1000 steps but uses an accelerated sampling strategy that reduces inference steps to 500.
Figure 11: Analysis of diffusion steps: Evaluation metrics include map generation IoU, time cost, and success rate on the Gibson validation navigation dataset.
4.4 Visualization of navigating with DAR.
@@
We follow the SGM configuration, and at the beginning of the navigation stage, the agent’s collected environmental information is insufficient to support accurate visual reasoning, so we use the frontier-based exploration strategy [28]. As more scenes are observed, the DAR model successfully predicts the position of the target, even though it has not yet been observed. During the navigation process, the DAR framework dynamically predicts object distributions in unexplored areas by leveraging the agent’s accumulated environmental memory. This predictive capability enables DAR to generate long-term goals that effectively guide the exploration trajectory, as shown in Fig. 12.
@@
Notably, the diffusion model’s inherent stochasticity introduces variability into DAR’s semantic reasoning. Consequently, even when operating in identical environments with the same initial conditions and target objectives, DAR exhibits non-deterministic behavior - each navigation query produces distinct exploratory paths and outcomes. As shown in Fig. 12, although the paths obtained are not completely identical, overall, DAR still significantly improves the efficiency of ObjectNav. From Fig. 9, it can also be seen that the DAR model not only successfully predicts the potential location of the target object during the navigation process, but also estimates the approximate locations of other objects and their context. There is still some deviation between the predicted target position and the true position by DAR, but we believe this is acceptable, as even humans cannot precisely locate unseen objects in an unseen scene.
@@
These two visualizations (Fig. 12 and 9) show that our DAR effectively expands limited local maps by generating unobserved regions, enabling the agent to infer the target’s location.
4.5 Ablation study.
Table 2: Performance of DAR model configurations on the Gibson and MP3D validation navigation datasets.
Model Gibson MP3D
SR(%$\uparrow$ SPL(%$\uparrow$ SR(%$\uparrow$ SPL(%$\uparrow$
Diff-M1 w/o RG 73.55 40.7 36.21 13.1
@@
Diff-M1 w/o SA 80.18 44.6 39.19 13.9
@@
Diff-M2 80.55 44.6 39.86 14.4
Diff-M1 81.23 45.3 41.44 15.2
In this section, we evaluate the performance of the three models proposed in Fig. 5 on the navigation dataset, measuring SR and SPL. The results are presented in Table. 2. Model 1 (M1) significantly outperforms the other two models. M3 employs the simplest room guidance fusion method but achieves the worst performance. Further ablation of M1 by removing the Semantic Attention (SA) module and Room Feature Fusion (RF) module leads to a performance drop of approximately 1%~2%.
As shown in the table, the use of room guidance leads to a significant performance improvement (accuracy metrics increase by 10%~15%). In the first row of the table ("w/o RG"), we disable all room guidance branches in Fig. 5, retaining only the time step $t$ and noise $x_{t}$ as inputs.
As shown in Tables. 1 and 2, if a map completion model achieves strong performance on our designed Map Completion Test Dataset, it also attains higher metrics on the navigation validation set. In conclusion, our experiments demonstrate the effectiveness of room guidance—leveraging LLM-inferred frontiers room categories to guide the diffusion model in generating object semantic maps for the corresponding regions significantly improves map completion accuracy and, consequently, Object Navigation success rates.
4.6 Comparisons with the related works.
Table 3: Performance comparison with other object navigation works.
ID Methods Gibson MP3D
SR(%$\uparrow$ SPL(%$\uparrow$ DTS(m)$\downarrow$ SR(%$\uparrow$ SPL(%$\uparrow$ DTS(m)$\downarrow$
I DD-PPO [45] 15.0 10.7 3.24 8.0 1.8 6.94
Red-Rabbit [48] - - - 34.6 7.9 -
THDA [29] - - - 28.4 11.0 5.62
SSCNav [25] - - - 27.1 11.2 5.71
Habitat-web [35] - - - 35.4 10.2 -
EmbCLIP [23] 68.1 39.5 1.15 29.2 10.1 5.40
ENTL [24] - - - 17.0 5.0 -
@@
ANS [6] 67.1 34.9 1.66 21.2 9.2 6.31
SemExp [5] 71.1 39.6 1.39 28.3 10.9 6.06
PONI [32] 73.6 41.0 1.25 27.8 12.1 5.63
L2M [18] - - - 32.1 11.0 5.12
Stubborn [28] - - - 31.2 13.5 5.01
L3MVN [50] 76.9 38.8 1.01 - - -
SGM [54] 78.0 44.0 1.11 37.7 14.7 4.93
T-Diff [51] 79.6 44.9 1.00 39.6 15.2 5.16
[…]
DAR(1K) 81.2 45.3 1.12 41.4 15.2 4.82
We categorize existing Object Navigation works, evaluated on either the Gibson or MP3D validation datasets, into two approaches: 1) End-to-End: Directly outputs actions, primarily using reinforcement learning (id I). 2) Modular-Based: Predicts a map-based long-term goal (id II). These are listed in Table 3. We briefly introduce the methods listed in the table as follows:
  • $\bullet$
@@
  • $\bullet$
    THDA: Introduces "Treasure Hunt Data Augmentation" to improve rewards and adaptability in new scenes.
  • $\bullet$
    SSCNav: an algorithm that explicitly models scene priors using a confidence-aware semantic scene completion module to complete the scene and guide the agent’s navigation
  • $\bullet$
@@
  • $\bullet$
    EmbCLIP: Integrates CLIP as a visual encoder into the model to predict the next action.
  • $\bullet$
    ENTL: A method for extracting long-sequence representations in embodied navigation, using vector-quantized predictions of future states conditioned on current states and actions.
  • $\bullet$
@@
  • $\bullet$
    L2M: actively learns to generate semantic maps outside the field of view of the agent and leverages the uncertainty over the semantic classes in the unobserved areas to decide on long term goals.
  • $\bullet$
@@
  • $\bullet$
    L3MVN: Leverages Large Language Models (LLMs) to incorporate commonsense reasoning for object searching, introducing two paradigms—zero-shot and feed-forward approaches—that use language to identify relevant frontiers from the semantic map as long-term goals.
  • $\bullet$
@@
  • $\bullet$
    T-Diff: Learns the distribution of trajectory sequences conditioned on current observations and goal states.
Our proposed DAR model, based on a map-generation approach for long-term goal reasoning, outperforms both map-completion methods (e.g., SGM, SSCNav, and L2M) and location-prediction methods (e.g., PONI, SemExp, and T-Diff). Notably, on the Gibson dataset, DAR surpasses the previous state-of-the-art (SOTA) supervised method, PONI, by 10.3% in Success Rate (SR), 10.5% in Success weighted by Path Length (SPL), and reduces Distance to Success (DTS) by 0.13 meters. Compared to the previous SOTA location-prediction method (T-Diff), DAR achieves nearly identical performance with 0.5K diffusion steps and further improves SR by 2%~4% with 1K steps.
Figure 12: Visualization of navigation using the proposed DAR method. The left four columns show the agent’s RGB view and local semantic map, including the trajectory and long-term goal (marked by a deep blue dot). It also displays the map generated by the DAR model. The right two columns show the path taken by the agent during navigation, demonstrating that the DAR method is more efficient than the PONI method.
5 Conclusions
In this work, we propose a modular-based self-supervised method for the object navigation task, named DAR. We demonstrate the effectiveness of applying diffusion models to lone-term goal reasoning.
We propose a Room Guidance Condition method integrated into the Denoising Diffusion Probabilistic Model (DDPM) framework for map generation and introduce a denoising model incorporating room guidance. Initially, we trained the DDPM on an indoor semantic map dataset, successfully generating high-quality and structurally plausible new map samples. This demonstrates that the trained DDPM effectively captures the statistical distribution patterns governing spatial relationships among indoor objects. Furthermore, we highlight that leveraging these patterns can significantly enhance the success rate and efficiency of agents in Object Navigation tasks.
Subsequently, we use Chain-of-Thought and positive-negative prompting techniques to query LLMs to obtain common sense knowledge, and incorporate it into the diffusion model’s reasoning process through the proposed Room Guidance method.
In summary, our DAR method explicitly leverages the structured knowledge from the Object-Room priors extracted from LLM and implicitly utilizes the unstructured knowledge from the pre-trained diffusion model. A series of experiments have shown that the DAR method achieves comparable performance to previous methods.
Acknowledge
This work was supported by by the Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 52105014)